# Language Mixing in Reasoning Language Models: Patterns, Impact, and Internal Causes

### v1

@@ Abstract
Mingyang Wang1,2,3 Lukas Lange1 Heike Adel4 Yunpu Ma2,3 Jannik Strötgen5 Hinrich Schütze2,3 1Bosch Center for Artificial Intelligence, Renningen, Germany 2LMU Munich, Germany 3Munich Center for Machine Learning (MCML) 4Hochschule der Medien, Stuttgart, Germany 5Karlsruhe University of Applied Sciences, Germany mingyang.wang2@de.bosch.com
@@ 2 Related Work
Reasoning language models, such as OpenAI’s o1/o3 (Jaech et al., 2024; OpenAI, 2025) and the DeepSeek-R1 series (Guo et al., 2025), have demonstrated strong capabilities on complex tasks by generating structured intermediate steps. While prior work has largely focused on improving reasoning quality via prompting or training, little attention has been paid to RLM behavior in multilingual settings. One underexplored aspect is language mixing, where the language used in reasoning steps differs from that of the input prompt. DeepSeek-R1 introduces a language consistency reward to reduce such mixing (Guo et al., 2025), but reports a drop in performance, suggesting potential benefits of language mixing. In contrast, Xie et al. (2025) show that responses without language mixing yield higher accuracy in logical reasoning tasks.
@@
A growing body of work has explored how multilingual models internally represent and process information across languages. Works by Wendler et al. (2024), Dumas et al. (2024), and Wang et al. (2025) reveal that models like Llama tend to rely on English representations internally, even when operating in other languages, highlighting a strong bias toward high-resource languages. While these works focus on internal activations using translation or knowledge probing tasks, our analysis links internal processing to external reasoning traces, showing that language mixing aligns with the model’s latent language preferences during reasoning.
@@
Prior work shows that current language models still struggle to understand and generate code-switched text, particularly in low-resource settings (Khanuja et al., 2020; Winata et al., 2023; Yong et al., 2023; Zhang et al., 2023; Li et al., 2024a, b). Recent studies have highlighted unnatural language confusion in multilingual models as another type of code-switching, such as source-language hallucination or off-target translation, especially in English-centric models under zero-shot conditions (Marchisio et al., 2024).
@@ 3 Experimental Setup
We evaluate a broad range of reasoning language models, including DeepSeek-R1 (Guo et al., 2025) and its distilled variants, DeepSeek-R1-Distill-Qwen-{1.5B, 7B, 14B, 32B} and DeepSeek-R1-Distill-Llama-{8B, 70B}. We also include QwQ-32B (QwenTeam, 2025b), Qwen3-{4B, 30B-A3B, 32B} (QwenTeam, 2025a), and Gemini 2.0 Flash Thinking (Google, 2025) models to broaden coverage. For comparing language mixing between RLMs and their backbones, we include Qwen2.5-{14B, 32B} and Llama3.3-70B-Instruct, to compare DeepSeek-R1-Distill-Qwen-{14B, 32B} and DeepSeek-R1-Distill-Llama-70B with their distillation backbones. Table 5 in the appendix summarizes all models used in our evaluation. We follow the official hyperparameter settings for each model (see Table 6 in Appendix A.1.1). Some reasoning models occasionally exhibit endless reasoning — continue to generate reasoning steps without reaching a final answer.444This overthinking behavior is also observed in prior work such as Cuadron et al. (2025). As this occurs only in specific cases for most models and does not reflect a consistent pattern, our analysis focuses on valid reasoning traces that conclude with a final answer.555See Appendix A.2.3 for validity statistics and accuracies.
@@ 5.1 Method: Script-controlled Generation
We constrain models to generate reasoning steps using only specific script(s) by masking the logits of all other scripts during decoding. While language-level control is difficult due to token overlap, script-level control offers a clean separation based on Unicode, as tokens from different scripts do not overlap. We apply this script control during the reasoning phase (though it could also be applied to the answer phase); the final answer is generated freely after the “</think>” token, as models typically default to the input language for answers (as shown in Section 4).
@@ 5.2 Results
For Arabic, Hindi, and Japanese, forcing reasoning in Latin or Han scripts significantly improves performance. In Hindi, for example, switching to Latin yields gains ranging from 44% to 115%. In contrast, using the native script consistently results in the lowest accuracy, indicating that models struggle with scripts underrepresented in their training data. For multi-script control (see Table 13 and Table 9 in the appendix), reasoning in both Latin and Han scripts does not outperform single-script control, and including the input script alongside Latin and/or Han leads to suboptimal results.
@@ 6.2 Results
As task difficulty increases, Latin usage rises while Devanagari decreases. The same trend appears in the reasoning traces (Figure 7(d)), suggesting a strong link between internal processing and external reasoning behavior. To quantify this connection, we calculate the Pearson correlation between the script usage across difficulty levels in hidden layers and reasoning traces, shown in Table 4. We observe consistently high correlations for both Arabic and Hindi inputs in R1-70B, R1-32B and R1-14B models. These results provide evidence that language mixing in reasoning traces reflects internal processing patterns. In particular, the internal preference for the Latin script explains why RLMs tend to mix into Latin-script reasoning, especially when processing underrepresented scripts.
@@ Acknowledgments
This work was partially supported by Deutsche Forschungsgemeinschaft (project SCHU 2246/14-1). We thank JUREAP 73 for supporting this work with computing resources. We also thank Sebastian Gerstner, Ahmad Dawar Hakimi, Lea Hirlimann, Valentin Knappich, Felicia Körner, Yihong Liu, Mohsen Mesgar, Ali Modarressi, Philipp Mondorf, Timo Pierre Schrader, Leonor Veloso, Wei Zhou, Yuqicheng Zhu for their valuable discussions.
@@ A.2.1 Language Mixing Patterns
In figs. 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21 and 22 and figs. 23, 24, 25, 26, 27, 28, 29, 30, 31, 32 and 33, we visualize the language composition of reasoning traces and final answers across models on the K&K and m-MMLU datasets, respectively. These figures provide a detailed view of how different RLMs vary in their use of languages during the reasoning process across inputs. Missing bars in some plots correspond to rare failure cases where the model enters an endless reasoning loop and fails to produce a final answer. For more details on these invalid generations, see Appendix A.2.3.
@@ A.2.2 Script-Controlled Generation
Tables 13 and 9 present the detailed performance results of three DeepSeek-R1 models (R1-70B, R1-32B, and R1-14B) across various script control strategies for languages from the K&K dataset. We compare model accuracy under three settings: no control (default decoding), single-script control (forcing reasoning in one script), and multi-script control (allowing reasoning in a combination of scripts). The languages are grouped into two sets based on their writing scripts: Arabic, Hindi, and Japanese use non-Latin/Han scripts; English, French, and Chinese use Latin or Han scripts. As shown, non-Latin/Han languages benefit substantially from being forced to reason in Latin or Han scripts, whereas native-script reasoning is optimal for Latin and Han-script languages. These findings highlight the importance of aligning reasoning scripts with the model’s internal preferences for achieving optimal multilingual reasoning performance.
@@ A.2.3 Model Performance Details
Here, we report the valid reasoning rate and accuracy across all reasoning models used in this work in tables 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 and 21. The valid reasoning rate refers to the proportion of generations that produce a complete reasoning trace followed by a final answer, as defined in Section 3. The accuracy is computed over all generations, both valid and invalid, and reflects the overall correctness of the final answers.

### v2

@@ Abstract
Mingyang Wang1,2,3 Lukas Lange1 Heike Adel4 Yunpu Ma2,3 Jannik Strötgen5 Hinrich Schütze2,3 1Bosch Center for Artificial Intelligence, Renningen, Germany 2LMU Munich, Germany 3Munich Center for Machine Learning (MCML) 4Hochschule der Medien, Stuttgart, Germany 5Karlsruhe University of Applied Sciences, Germany mingyang@cis.lmu.de
@@ 2 Related Work
Reasoning language models, such as OpenAI’s o1/o3 (Jaech et al., 2024; OpenAI, 2025) and the DeepSeek-R1 series (Guo et al., 2025), have demonstrated strong capabilities on complex tasks by generating structured intermediate steps (Chen et al., 2025). While prior work has largely focused on improving reasoning quality via prompting or training, little attention has been paid to RLM behavior in multilingual settings. One underexplored aspect is language mixing, where the language used in reasoning steps differs from that of the input prompt. DeepSeek-R1 introduces a language consistency reward to reduce such mixing (Guo et al., 2025), but reports a drop in performance, suggesting potential benefits of language mixing. In contrast, Xie et al. (2025) show that responses without language mixing yield higher accuracy in logical reasoning tasks.
@@
A growing body of work has explored how multilingual models internally represent and process information across languages. Works by Wendler et al. (2024), Dumas et al. (2024), Wang et al. (2025), and Liu et al. (2025) reveal that models like Llama tend to rely on English representations internally, even when operating in other languages, highlighting a strong bias toward high-resource languages. While these works focus on internal activations using translation or knowledge probing tasks, our analysis links internal processing to external reasoning traces, showing that language mixing aligns with the model’s latent language preferences during reasoning.
@@
Prior work shows that current language models still struggle to understand and generate code-switched text, particularly in low-resource settings (Khanuja et al., 2020; Winata et al., 2023; Yong et al., 2023; Zhang et al., 2023; Li et al., 2024a, b). Recent studies have highlighted unnatural language confusion in multilingual models as another type of code-switching (Marchisio et al., 2024; Nie et al., 2025), such as source-language hallucination or off-target translation, especially in English-centric models under zero-shot conditions.
@@ 3 Experimental Setup
We evaluate a broad range of reasoning language models, including DeepSeek-R1 (Guo et al., 2025) and its distilled variants, DeepSeek-R1-Distill-Qwen-{1.5B, 7B, 14B, 32B} and DeepSeek-R1-Distill-Llama-{8B, 70B}. We also include QwQ-32B (QwenTeam, 2025b), Qwen3-{4B, 30B-A3B, 32B} (QwenTeam, 2025a), and Gemini 2.0 Flash Thinking (Google, 2025) models to broaden coverage. For comparing language mixing between RLMs and their backbones, we include Qwen2.5-{14B, 32B} and Llama3.3-70B-Instruct, to compare DeepSeek-R1-Distill-Qwen-{14B, 32B} and DeepSeek-R1-Distill-Llama-70B with their distillation backbones. Table 5 in Appendix A.1.1 summarizes all models used in our evaluation. We follow the official hyperparameter settings for each model (see Table 6 in Appendix A.1.1). Some reasoning models occasionally exhibit endless reasoning — continue to generate reasoning steps without reaching a final answer.444This overthinking behavior is also observed in prior work such as Cuadron et al. (2025). As this occurs only in specific cases for most models and does not reflect a consistent pattern, our analysis focuses on valid reasoning traces that conclude with a final answer.555See Appendix A.2.3 for validity statistics and accuracies.
@@ 5.1 Method: Script-controlled Generation
We constrain models to generate reasoning steps using only specific script(s) by masking the logits of all other scripts during decoding. While language-level control is difficult due to token overlap, script-level control offers a clean separation based on Unicode, as tokens from different scripts do not overlap. We apply this script control during the reasoning phase (though it could also be applied to the answer phase); the final answer is generated freely after the “</think>” token, as models typically default to the input language for answers (as shown in Section 4). For examples of model generation under different script control settings, see Figures˜11, 12, 13 and 14 in Appendix A.2.2.
@@ 5.2 Results
For Arabic, Hindi, and Japanese, forcing reasoning in Latin or Han scripts significantly improves performance. In Hindi, for example, switching to Latin yields gains ranging from 44% to 115%. In contrast, using the native script consistently results in the lowest accuracy, indicating that models struggle with scripts underrepresented in their training data. For multi-script control (see Table 13 and Table 11 in Appendix A.2.2), reasoning in both Latin and Han scripts does not outperform single-script control, and including the input script alongside Latin and/or Han leads to suboptimal results.
@@ 6.1 Method: Logit Lens Analysis
To quantify this connection, we compute the Pearson correlation between script usage across difficulty levels in hidden layers and reasoning traces. Specifically: (1) for each difficulty level (ranging from 2 to 8 ppl), we calculate the proportion of tokens belonging to different scripts (e.g., Latin, Devanagari) in both the latent space (estimated with the Logit Lens) and the reasoning trace (measured via script-level detection); and (2) for each script, we compute the Pearson correlation coefficient between its percentage across difficulty levels in the latent space and its corresponding percentage in the reasoning trace. A more detailed description of the Pearson correlation calculation is provided in Appendix A.1.3.
@@ 6.2 Results
As task difficulty increases, Latin usage rises while Devanagari decreases. The same trend appears in the reasoning traces (Figure 7(d)), suggesting a strong link between internal processing and external reasoning behavior. The Pearson correlations between script usage across difficulty levels in hidden layers and reasoning traces are reported in Table 4. We observe consistently high correlations for both Arabic and Hindi inputs in R1-70B, R1-32B and R1-14B models. These results provide evidence that language mixing in reasoning traces reflects internal processing patterns. In particular, the internal preference for the Latin script explains why RLMs tend to mix into Latin-script reasoning, especially when processing underrepresented scripts.
@@ 7 Discussion and Future Works
Adaptive reasoning language selection.
Our results show that switching to high-resource languages during reasoning can substantially improve model performance on under-resourced inputs. This suggests a promising direction for future work: training RLMs to adaptively determine which language to use at different stages of reasoning, rather than relying on static language constraints. Such adaptive control could allow the model to dynamically leverage the strengths of high-resource languages while still accommodating the input language, ultimately realizing more robust multilingual reasoning. This idea is aligned with recent advances in adaptive memory usage for reasoning (Yan et al., 2025), which highlight the benefits of enabling models to flexibly select internal strategies based on task demands.
@@ Acknowledgments
This work was partially supported by Deutsche Forschungsgemeinschaft (project SCHU 2246/14-1) and the EU Project SMARTY (GA 101140087). We thank JUREAP 73 for supporting this work with computing resources. We also thank Sebastian Gerstner, Ahmad Dawar Hakimi, Lea Hirlimann, Valentin Knappich, Felicia Körner, Yihong Liu, Ali Modarressi, Philipp Mondorf, Timo Pierre Schrader, Leonor Veloso, Wei Zhou, Yuqicheng Zhu for their valuable discussions.
@@ A.1.2 Datasets
Language | Abbreviation | Script |
Languages of investigation | ||
Arabic | ar / AR-XY | Arabic |
Bengali | BN-BD | Bangla |
Chinese | zh / ZH-CN | Han |
English | en / EN-US | Latin |
French | fr | Latin |
German | DE-DE | Latin |
Hindi | hi / HI-IN | Devanagari |
Indonesian | ID-ID | Latin |
Italian | IT-IT | Latin |
Japanese | ja / JA-JP | Han, Hiragana, Katakana |
Korean | KO-KR | Hangul |
Portuguese | PT-BR | Latin |
Spanish | ES-LA | Latin |
Swahili | SW-KE | Latin |
Yoruba | YO-NG | Latin |
Languages emerge in the reasoning trace | ||
Persian | fa | Perso-Arabic |
Marathi | mr | Devanagari |
Russian | ru | Cyrillic |
@@
A.1.3 Pearson correlation calculation
To clarify how the Pearson correlation coefficients in Table 4 are calculated, we provide a detailed description of the procedure.
We take Figure 7 in the main text as an example (which shows the script composition of internal representations and reasoning traces for Hindi inputs in the R1-70B model). The Pearson correlation is computed as follows:
1.
For each difficulty level (from 2 to 8 ppl), we collect the percentage of tokens belonging to different scripts, such as Latin and Devanagari. These percentages are measured in (1) the latent space, obtained via Logit Lens analysis of hidden layer activations; and (2) the reasoning trace, obtained through script-level detection of the generated text.
2.
For each script, we then calculate the Pearson correlation coefficient between its sequence of percentages across difficulty levels in the latent space and the corresponding sequence in the reasoning trace. An illustrative example is provided in Table 9.
These Pearson correlations indicate how closely the model’s internal script usage aligns with the script composition of the generated reasoning trace.
Difficulty | Latin – latent (%) | Latin – reasoning (%) | Devanagari – latent (%) | Devanagari – reasoning (%) |
2 ppl | 77.01 | 0.76 | 17.82 | 99.19 |
3 ppl | 77.00 | 4.22 | 18.34 | 95.75 |
4 ppl | 78.54 | 15.05 | 16.96 | 78.11 |
5 ppl | 80.93 | 31.68 | 15.11 | 61.90 |
6 ppl | 84.78 | 48.02 | 12.18 | 48.02 |
7 ppl | 91.46 | 44.71 | 6.60 | 44.33 |
8 ppl | 91.79 | 63.26 | 6.40 | 21.54 |
@@ A.2.1 Language Mixing Patterns
In Figures˜15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25 and 26 and Figures˜27, 28, 29, 30, 31, 32, 33, 34, 35, 36 and 37, we visualize the language composition of reasoning traces and final answers across models on the K&K and m-MMLU datasets, respectively. These figures provide a detailed view of how different RLMs vary in their use of languages during the reasoning process across inputs. Missing bars in some plots correspond to rare failure cases where the model enters an endless reasoning loop and fails to produce a final answer. For more details on these invalid generations, see Appendix A.2.3.
@@ A.2.2 Script-Controlled Generation
Script control complete results.
Tables 13 and 11 present the detailed performance results of three DeepSeek-R1 models (R1-70B, R1-32B, and R1-14B) across various script control strategies for languages from the K&K dataset. We compare model accuracy under three settings: no control (default decoding), single-script control (forcing reasoning in one script), and multi-script control (allowing reasoning in a combination of scripts). The languages are grouped into two sets based on their writing scripts: Arabic, Hindi, and Japanese use non-Latin/Han scripts; English, French, and Chinese use Latin or Han scripts. As shown, non-Latin/Han languages benefit substantially from being forced to reason in Latin or Han scripts, whereas native-script reasoning is optimal for Latin and Han-script languages. These findings highlight the importance of aligning reasoning scripts with the model’s internal preferences for achieving optimal multilingual reasoning performance.
Script control generation examples.
We demonstrate the effects of script control on multilingual reasoning through examples using the same Knights and Knaves puzzle with Arabic input (Figures˜11, 12, 13 and 14). Figure 11 shows unconstrained reasoning with natural language mixing across English, Chinese, and Arabic. Figures 12, 13, and 14 demonstrate controlled reasoning under Latin, Han, and Arabic script constraints respectively. These examples illustrate how script control can effectively guide language selection while maintaining the model’s cross-lingual reasoning capabilities.
@@ A.2.3 Model Performance Details
Here, we report the valid reasoning rate and accuracy across all reasoning models used in this work in Tables˜12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22 and 23. The valid reasoning rate refers to the proportion of generations that produce a complete reasoning trace followed by a final answer, as defined in Section 3. The accuracy is computed over all generations, both valid and invalid, and reflects the overall correctness of the final answers.